OpenStudy (anonymous):

rationalize the denominator of 2/∛5

OpenStudy (anonymous):

\[2/\sqrt[3]{5}=2\sqrt[3]{25}/\sqrt[3]{5}\sqrt[3]{25}=2\sqrt[3]{25}/5\]

OpenStudy (anonymous):

On a square root, you multiply it by the same factor to rationalize. On a cube root, the number of factors needs to be three, so we multiply by the cube root squared, or cube root of the square.

OpenStudy (anonymous):